including 6.1% sales tax, a diamond ring sold for $2864.70. Find the price of the ring before the tax was added. Round to nearest cent.
Answer by JulietG(1812)

100% of the price + 6.1% of the price (that's a very low sales tax rate!) = $2864.70
106.1 * x = 2864.70
Divide each side by 106.1 (and then multiply by 100 -- because we used 6.1 rather than .061)
$2700
.
You can prove this by multiplying the sales tax of .061 * 2700 = 164.70. 2700 + 164.70 = 2864.70